Help
6. A 2370 kg airplane starts from rest and, with a constant acceleration of 5.0 m/s, reaches its takeoff speed in 9.00 s. a. What is the net force on the airplane during this time? b. What is the airplanes takeoff speed? c. How far down the runway does the airplane travel before taking off? 7. A train engine pulls a freight car that has a mass of 5.0 x 10 kg with a constant force. The freight car at rest, then travels 43 m on the straight track in the next 20.0 s. What is the net force on the starts freight car? A model sailboat is slowly sailing east across a pond at 0.33 m/s. A gust of wind gives the sailboat a constant acceleration of 0.30 m/s (towards the east) during a time interval of 2.0 s. During this time the net force on the sailboat is 0.375 N a. What is the mass of the sailboat? b. How far does the sailboat travel during the 2.0 s time interval? c. What is the sailboats velocity at the end of the 2.0 s time interval? 8.
Answer #1

6] a] Net force on the airplane = F = ma = 2370 x 5 = 11850 N

b] Airplane's takeoff speed will be:

v = u + at = 0 + 5(9) = 45 m/s

c] Distance the airplane travels before taking off will be:

S = ut + (1/2)at2 = 0 + (1/2)(5)(9)2 = 202.5 meters.